# Are Large Language Models Good Statisticians?

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities across a range of scientific tasks including mathematics, physics, and chemistry. Despite their successes, the effectiveness of LLMs in handling complex statistical tasks remains systematically under-explored. To bridge this gap, we introduce StatQA, a new benchmark designed for statistical analysis tasks. StatQA comprises 11,623 examples tailored to evaluate LLMs’ proficiency in specialized statistical tasks and their applicability assessment capabilities, particularly for hypothesis testing methods. We systematically experiment with representative LLMs using various prompting strategies and show that even state-of-the-art models such as GPT-4o achieve a best performance of only 64.83%, indicating significant room for improvement. Notably, while open-source LLMs (e.g., LLaMA-3) show limited capability, those fine-tuned ones exhibit marked improvements, outperforming all in-context learning-based methods (e.g., GPT-4o). Moreover, our comparative human experiments highlight a striking contrast in error types between LLMs and humans: LLMs primarily make applicability errors, whereas humans mostly make statistical task confusion errors. This divergence highlights distinct areas of proficiency and deficiency, suggesting that combining LLM and human expertise could lead to complementary strengths, inviting further investigation into their collaborative potential.
@@ 1 Introduction
Statistical analysis can capture data patterns and convert them into usable evidence, which is crucial to data science and machine learning applications [1, 2, 3]. As shown in Figure 1, a typical statistical analysis task involves, given a table $D$ and a statistical question $Q$, a qualified statistician should be proficient in selecting relevant columns $\mathbb{C}$, choosing the appropriate statistical methods $\mathbb{M}$, and computing the results based on $\mathbb{M}$ using $\mathbb{C}$. Formally, the statistical analysis task can be divided into two stages: (1) identifying appropriate statistical methods and parameters (e.g., relevant columns in the table), and (2) computing the statistical results and deriving the conclusion. The first stage requires statistical expertise to assess the applicability of methods, considering factors such as data type, distribution characteristics, and sample size, which is the core of the statistical task. In contrast, the second stage (i.e., computation) can be easily aided by external tools [4] such as WolframAlpha [5].
@@
Inspired by the extensive application of Large Language Models (LLMs) [6], we pose a critical question: Do LLMs truly understand such “statistical literacy”? Specifically, can LLMs competently select relevant data, recognize prerequisites, and discern appropriate usage scenarios to assess the effectiveness of statistical methods?
@@
    Our source code and data are available at https://github.com/derrickzhuyz/StatQA
@@ 2.3 StatQA Construction
We collect 78 tables from real-world applications, covering various domains including education, medicine, engineering, and the economy. These tables are carefully gathered by post-graduate students in statistics from Kaggle [14] and Rdatasets [15]. The metadata includes descriptive information about the tabular data and details of each data column. Specifically, it encompasses the column header, data type, normality, and description, which are crucial for prerequisite checks in statistical analysis. Descriptive information in the metadata can be obtained from their sources, while the data types and normality of each column can be calculated and derived from the tabular data. For tables lacking metadata, we performed manual annotations. For tables with existing metadata, we conducted manual validation to ensure accuracy.
@@
For a statistical question, if the prerequisites for the methods are not met, we label it as “hard” due to potential confusion; otherwise, it is labeled as “easy”. Based on these labels, we expand the underrepresented categories and sample the abundant ones to obtain a balanced synthesized examples in terms of task difficulty. Next, we split these synthesized examples to ensure no table overlaps in training and test sets: data synthesized from source tables No. 1 to 36 will be used for subsequent tests (i.e., StatQA), while tables No. 37 to 78 will be used for training (i.e., training set $\mathbb{D}_{\text{train}}$). The $\mathbb{D}_{\text{train}}$ is used to fine-tuning LLMs.
@@
Inspired by LLMs’ capabilities in text comprehension and processing, we use GPT-3.5-Turbo to refine the phrasing and expression of statistical questions for the test data. We provide GPT-3.5-Turbo with origin questions and their descriptive information, instructing GPT-3.5-Turbo to paraphrase and refine the question sentences without changing the original meaning, aiming for more coherent and diverse expressions, which is noted as $Q^{*}(\mathbb{C})$.
Discussion of Quality Control. We devise several strategies to ensure the high quality of our StatQA. (1) Quality of Question Templates: The quality of question templates is a focal point in our quality control. We categorize templates based on statistical tasks and applicable scenarios of methods. For each category, we meticulously prepare 10 to 20 templates for random selection to enhance diversity. To ensure the templates are representative, we recruit two post-graduate students in statistics to design and review them. (2) Question Refinement: To improve the coherence and flexibility of the questions generated by the templates, we conduct question refinement using GPT-3.5-Turbo. Note that question refinement is exclusively performed on StatQA to increase diversity and ensure differences from the training set. (3) Expert Reviews: Last but most importantly, we conduct manual reviews by two post-graduate students in statistics to carefully check all examples in StatQA.
@@ 3.1 Setup
We design experiments for LLMs similar to human statisticians’ mindset, as presented in Figure 1, to evaluate the abilities of LLMs in statistical tasks. Because of the limitation of input tokens, we provide LLMs column information instead of the whole table, including the column headers, number of rows, data type, and normality. In the experiment, the LLMs need to pick headers of relevant data columns, assess the methods’ applicability, and select all statistical methods that fit the usage scenario and prerequisites as statisticians, then respond in a specific format. Since LLM responses might be invalid or include irrelevant content, cleaning and extraction are necessary, then comparing the extracted answers to the ground truth for evaluation. In the human experiments, we use the same protocol for consistency and develop a testing platform to facilitate participant selection. More details of experimental setups including hyperparameters, prompts, procedures and GUI of the testing platform used in human experiments, are provided in Section C.3.
@@
Accuracy of relevant data columns and applicable methods selections, noted as $Acc(\mathbb{C},\mathbb{M})$, is used as our metrics to evaluate if LLMs or participants truly understand the question and the applicability of statistical methods. $Acc(\mathbb{C},\mathbb{M})$ refers to the proportion of methods and column selections fully aligned with the ground truth without any omissions or incorrect selections.
@@ 3.1.2 Human Experiments
Table 2: Experimental Results of $Acc(\mathbb{C},\mathbb{M})$(%) on mini-StatQA. The 1st place, 2nd place, 3rd place results in all experiments are highlighted in red, blue, and green respectively. The bold results are the best in each section. The underlined results are the leading ones of each subgroup in overall $Acc(\mathbb{C},\mathbb{M})$, whose performances in subtasks are shown in Figure 5.
@@ 3.2 Experimental Results and Analysis
Figure 5: Best Results of Each Model in Sub-tasks.
@@
We observe that compared to 0-shot, 1-shot learning increased the $Acc(\mathbb{C},\mathbb{M})$ for the LLaMA-2/3 models by an average of 9.6%, and by 4.7% for GPT models. However, CoT leads to a minor performance decrease in most cases of LLaMA-2/3, whereas its impact on larger GPT models is negligible. More analysis for CoT is shown in Section D.2. The introduced domain knowledge prompting has proven effective, for six out of the seven non-fine-tuned LLMs obtained their best results in our experiments. Notably, it has a more pronounced effect on larger LLMs with stronger comprehension abilities, particularly GPT-4o.
@@ 3.3 Errors Analysis
We categorize errors into four distinct types and mixed errors, with examples in Section D.1. The error categories are (1) Invalid Answer: meaningless responses or do not conform to the required format; (2) Column Selection Error: irrelevant or incorrect column selection; (3) Statistical Task Confusion: confusion regarding the category of statistical tasks, leading to incorrect selection of methods; (4) Applicability Error: no confusion on task category but failing to discern the usage scenarios and prerequisites, resulting in the selection of inapplicable methods; (5) Mixed Errors: valid answer but contain multiple types of errors.
@@
Finding 5. LLMs are good at distinguishing different statistical tasks then selecting associated methods but struggle to utilize the domain knowledge to assess method applicability effectively. Conversely, humans excel at discerning method applicability but are prone to task confusion.
@@ 4 Research Opportunity
We establish StatQA, focusing on the evaluation of applicability rather than computational results, aiming to highlight the importance of discerning statistical method suitability. StatQA can be expanded to encompass a broader range of statistical tasks and methods, as discussed in Section F in the Appendix. Beyond the field of statistics, the evaluation of method applicability is also crucial in professional domains such as finance and operations research. Consequently, curating a more extensive benchmark represents a significant undertaking.
Human-AI Collaboration in the Statistical Task.
LLMs and humans exhibit distinct aspects of superior capability in statistical tasks. Therefore, an in-depth study into harnessing their unique strengths for complementary collaboration to achieve optimal performance can be a valuable endeavor. This approach could leverage the computational efficiency and data handling capabilities of LLMs alongside the nuanced understanding and domain knowledge of human experts, leading to more robust and accurate statistical analyses.
@@ 5 Related Work
Proprietary LLMs like ChatGPT [18] and GPT-4 [19] exhibit impressive capabilities in text comprehension and processing, and recent-released GPT-4o [20] is further elevated performance and speed. Open-source models are also favored for their flexibility and suitability for customization, among the most representative examples are the LLaMA series models, such as LLaMA-2 [16] and the new LLaMA-3 [17]. Meanwhile, prompting strategies [23] and fine-tuning methods like LoRA [22], enable LLMs to better adapt to specific tasks with less computational costs.
@@ 6 Conclusion
D: More Details of Error Analysis
E: Qualitative Analysis
F: Limitations
G: Ethic Statement
H: Datasheet for StatQA
@@ A.1 Large Language Models (LLMs)
LLMs are increasingly employed in both academic and industrial sectors, with extensive applications across disciplines such as science, education, medicine, and finance [33, 34, 35, 36, 37, 38]. Particularly since the advent of ChatGPT [18], LLMs have demonstrated impressive capabilities and versatility in text comprehension, question answering, and reasoning. The GPT-series model released by OpenAI is one of the most representative and widely used proprietary LLMs, including the newly released GPT-4o as well as ChatGPT and GPT-4, which have achieved state-of-the-art performance on numerous professional benchmarks without fine-tuning [18, 19, 20, 39]. Other renowned proprietary LLMs include Claude [40], Qwen [41], and Gemini [42].
@@ B.1 Access to StatQA
StatQA contains 11,623 examples and mini-StatQA contains 1,163 examples across five typical categories of statistical tasks. To facilitate further research, we make our code and data available at https://github.com/derrickzhuyz/StatQA, under GPL-3 license.
@@ C.3 Prompts
Appendix D More Details of Error Analysis
D.1 Examples of Error Taxonomy

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities across a range of scientific tasks including mathematics, physics, and chemistry. Despite their successes, the effectiveness of LLMs in handling complex statistical tasks remains systematically under-explored. To bridge this gap, we introduce StatQA, a new benchmark designed for statistical analysis tasks. StatQA comprises 11,623 examples tailored to evaluate LLMs’ proficiency in specialized statistical tasks and their applicability assessment capabilities, particularly for hypothesis testing methods. We systematically experiment with representative LLMs using various prompting strategies and show that even state-of-the-art models such as GPT-4o achieve a best performance of only 64.83%, indicating significant room for improvement. Notably, while open-source LLMs (e.g., LLaMA-3) show limited capability, those fine-tuned ones exhibit marked improvements, outperforming all in-context learning-based methods (e.g., GPT-4o). Moreover, our comparative human experiments highlight a striking contrast in error types between LLMs and humans: LLMs primarily make applicability errors, whereas humans mostly make statistical task confusion errors. This divergence highlights distinct areas of proficiency and deficiency, suggesting that combining LLM and human expertise could lead to complementary strengths, inviting further investigation into their collaborative potential. Our source code and data are available at https://statqa.github.io/.
@@ 1 Introduction
Statistical analysis can capture data patterns and convert them into usable evidence, which is crucial to data science and machine learning applications [1, 2, 3, 4, 5, 6, 7, 8]. As shown in Figure 1, a typical statistical analysis task involves, given a table $D$ and a statistical question $Q$, a qualified statistician should be proficient in selecting relevant columns $\mathbb{C}$, choosing the appropriate statistical methods $\mathbb{M}$, and computing the results based on $\mathbb{M}$ using $\mathbb{C}$. Formally, the statistical analysis task can be divided into two stages: (1) identifying appropriate statistical methods and parameters (e.g., relevant columns in the table), and (2) computing the statistical results and deriving the conclusion. The first stage requires statistical expertise to assess the applicability of methods, considering factors such as data type, distribution characteristics, and sample size, which is the core of the statistical task. In contrast, the second stage (i.e., computation) can be easily aided by external tools [9] such as WolframAlpha [10].
@@
Inspired by the extensive application of Large Language Models (LLMs) [11, 12, 13], we pose a critical question: Do LLMs truly understand such “statistical literacy”? Specifically, can LLMs competently select relevant data, recognize prerequisites, and discern appropriate usage scenarios to assess the effectiveness of statistical methods?
@@ 2.3 StatQA Construction
We collect 78 tables from real-world applications, covering various domains including education, medicine, economy, etc., as shown in Section B.5. These tables are carefully gathered by post-graduate students in statistics from Kaggle [21] and Rdatasets [22]. The metadata includes descriptive information about the tabular data and details of each data column. Specifically, it encompasses the column header, data type, normality, and description, which are crucial for prerequisite checks in statistical analysis. Descriptive information in metadata can be obtained from their sources, while the data types and normality of each column can be calculated and derived from the tabular data. For tables lacking metadata, we perform manual annotations. For tables with existing metadata, we conduct manual validation to ensure accuracy.
@@
For a statistical question, if the set of applicable methods constitutes a proper subset of the candidate methods (i.e., all methods for comparable scenarios), it indicates that certain prerequisites for some methods are not satisfied. In such cases, we label the question as “hard”, reflecting the increased challenge of accurately assessing applicability to eliminate inapplicable methods, otherwise the question is labeled as “easy”. Based on these labels, we expand the underrepresented categories and sample the abundant ones to obtain balanced synthesized examples in terms of task difficulty. Next, we split these synthesized examples to ensure no table overlaps in training and test sets: data synthesized from source tables No. 1 to 36 will be used for subsequent tests (i.e., StatQA), while tables No. 37 to 78 will be used for training (i.e., training set $\mathbb{D}_{\text{train}}$). The $\mathbb{D}_{\text{train}}$ is used to fine-tuning LLMs.
@@
Inspired by LLMs’ capabilities in text comprehension and processing, we use GPT-3.5-Turbo to refine the phrasing and expression of statistical questions for the test data. We provide GPT-3.5-Turbo with original questions and their descriptive information, instructing GPT-3.5-Turbo to paraphrase and refine the question sentences without changing the meaning, aiming for more coherent and diverse expressions, which is noted as $Q^{*}(\mathbb{C})$.
Discussion of Quality Control. We devise several strategies to ensure the high quality of our StatQA. (1) Quality of Question Templates: The quality of question templates is a focal point in our quality control. We categorize templates based on statistical tasks and applicable scenarios of methods. For each category, we meticulously prepare 10 to 20 templates for random selection to enhance diversity. To ensure the templates are representative, we recruit two post-graduate students in statistics to design and review them. (2) Question Refinement: The objectives of refinement are to correct potential grammar mistakes, improve semantical coherence, and increase the diversity of expressions for the questions generated by the templates. In practice, GPT-3.5-Turbo demonstrates a satisfactory level of English proficiency in achieving these goals and offers better cost-effectiveness; therefore, we employ GPT-3.5-Turbo as the refiner. Note that question refinement is exclusively performed on StatQA to increase diversity and ensure differences from the training set. The average BLEU [23], BERTScore [24] calculated between original questions and refined versions in StatQA is 0.126, 0.920 respectively. The low BLEU reflects notable vocabulary differences, while the high BERTScore suggests semantic consistency, collectively indicating adequate rephrasing that preserves the original meaning. (3) Expert Reviews: Last but most importantly, we conduct manual reviews by two post-graduate students in statistics to carefully check all examples in StatQA.
@@ 3.1 Setup
We design experiments for LLMs similar to human statisticians’ mindset, as presented in Figure 1, to evaluate the abilities of LLMs in statistical tasks. Because of the limitation of input tokens, we provide LLMs column information instead of the whole table, including the column headers, number of rows, data type, and normality. In the experiment, the LLMs need to pick headers of relevant data columns, assess the methods’ applicability, and select all statistical methods that fit the usage scenario and prerequisites as statisticians, then respond in a specific format. Since LLM responses might be invalid or include irrelevant content, cleaning and extraction are necessary, then comparing the extracted answers to the ground truth for evaluation. In the human experiments, we use the same protocol for consistency and develop a testing platform to facilitate participant selection. More details of experimental setups including hyperparameters, prompts, procedures and GUI of the testing platform used in human experiments, are provided in Section C.
@@
Accuracy of relevant data columns and applicable methods selections, noted as $Acc(\mathbb{C},\mathbb{M})$, is used as our metrics to evaluate if LLMs or participants truly understand the question and the applicability of statistical methods. $Acc(\mathbb{C},\mathbb{M})$ refers to the proportion of methods and column selections fully aligned with the ground truth without any omissions or incorrect selections:
$Acc(\mathbb{C},\mathbb{M})=\frac{\sum_{i=1}^{N}1(\hat{\mathbb{C}}_{i}=\mathbb{C}_{i},\hat{\mathbb{M}}_{i}=\mathbb{M}_{i})}{N}\text{,}$ (1)
where $N$ is the total number of test examples.
@@ 3.1.2 Human Experiments
Table 2: Experimental Results of $Acc(\mathbb{C},\mathbb{M})$(%) on mini-StatQA. The 1st, 2nd, 3rd place results in all experiments are highlighted in red, blue, and green respectively. The bold results are the best in each section. The underlined results are the leading ones of each subgroup in overall $Acc(\mathbb{C},\mathbb{M})$, whose performances in sub-tasks are shown in Figure 5. CA: Correlation Analysis; CTT: Contingency Table Test; DCT: Distribution Compliance Test; VT: Variance Test; DS: Descriptive Statistics.
@@ 3.2 Experimental Results and Analysis
Figure 5: Best Results of Each Model in Five Sub-tasks (Section 2.1).
@@
We observe that compared to 0-shot, 1-shot learning increased the $Acc(\mathbb{C},\mathbb{M})$ for the LLaMA-2/3 models by an average of 9.6%, and by 4.7% for GPT models. However, CoT leads to a minor performance decrease in most cases of LLaMA-2/3, whereas its impact on larger GPT models is negligible. More analysis for CoT is shown in Section E.2. The introduced domain knowledge prompting has proven effective, for six out of the seven non-fine-tuned LLMs obtained their best results in our experiments. Notably, it has a more pronounced effect on larger LLMs with stronger comprehension abilities, particularly GPT-4o.
@@ 3.3 Errors Analysis
We categorize errors into four distinct types and mixed errors, with examples in Section E.1. The error categories are (1) Invalid Answer: meaningless responses or do not conform to the required format; (2) Column Selection Error: irrelevant or incorrect column selection; (3) Statistical Task Confusion: confusion regarding the category of statistical tasks, leading to incorrect selection of methods; (4) Applicability Error: no confusion on task category but failing to discern the usage scenarios and prerequisites, resulting in the selection of inapplicable methods; (5) Mixed Errors: valid answer but contain multiple types of errors.
@@
Finding 5. LLMs are good at distinguishing different statistical tasks then selecting associated methods, but they struggle to utilize the domain knowledge to assess method applicability effectively. Conversely, humans excel at discerning method applicability but are prone to task confusion.
@@ 4 Research Opportunity
Human-AI Collaboration in the Statistical Task.
LLMs and humans exhibit distinct aspects of superior capability in statistical tasks. Therefore, an in-depth study into harnessing their unique strengths for complementary collaboration to achieve optimal performance can be a valuable endeavor. This approach could leverage the computational efficiency and data handling capabilities of LLMs alongside the nuanced understanding and domain knowledge of human experts, leading to more robust and accurate statistical analysis.
@@
We establish StatQA, focusing on the evaluation of applicability rather than computational results, aiming to highlight the importance of discerning statistical method suitability. StatQA can be expanded to encompass a broader range of statistical tasks and methods, as discussed in Section G in the Appendix. Beyond the field of statistics, the evaluation of method applicability is also crucial in professional domains such as finance and operations research. Consequently, curating a more extensive benchmark represents a significant undertaking.
@@ 5 Related Work
Proprietary LLMs like ChatGPT [27] and GPT-4 [28] exhibit impressive capabilities in text comprehension and processing, and recent-released GPT-4o [29] is further elevated performance and speed. Open-source models are also favored for their flexibility and suitability for customization, among the most representative examples are the LLaMA series models, such as LLaMA-2 [25] and the new LLaMA-3 [26]. Meanwhile, prompting strategies [32, 33] and fine-tuning methods like LoRA [31], enable LLMs to better adapt to specific tasks with less computational costs.
@@ 6 Conclusion
7 Acknowledgement
This paper is supported by NSF of China (62402409), Guangdong Basic and Applied Basic Research Foundation (2023A1515110545), CCF-Huawei Populus Grove Fund (CCF-HuaweiDB202403), and the Red Bird Grant, Research Grant from The Hong Kong University of Science and Technology (Guangzhou).
@@ 7 Acknowledgement
D: More Experimental Results
E: More Error Analysis
F: Qualitative Analysis
G: Limitations
H: Ethic Statement
I: Datasheet for StatQA
@@ A.1 Large Language Models (LLMs)
LLMs are increasingly employed in both academic and industrial sectors, with extensive applications across disciplines such as science, education, medicine, and finance [43, 44, 45, 46, 47, 48, 49]. Particularly since the advent of ChatGPT [27], LLMs have demonstrated impressive capabilities and versatility in data analysis, question answering, and reasoning [50, 51, 52, 53, 54, 55, 56, 57, 58]. The GPT-series model released by OpenAI is one of the most representative and widely used proprietary LLMs, including the newly released GPT-4o as well as ChatGPT and GPT-4, which have achieved state-of-the-art performance on numerous professional benchmarks without fine-tuning [27, 28, 29, 59]. Other renowned proprietary LLMs include Claude [60], Qwen [61], and Gemini [62].
@@ B.1 Access to StatQA
StatQA contains 11,623 examples and mini-StatQA contains 1,163 examples across five typical categories of statistical tasks. To facilitate further research, we make our code and data available at https://statqa.github.io/, under GPL-3 license.
@@ B.4 More Examples in StatQA
B.5 Domains of Source Tables
As mentioned, we have assembled 78 tables to serve as the foundation for the construction of StatQA, encompassing six domains in real-world scenarios: education, medicine, science, engineering, economy, and life. In the process of dataset partitioning (Step 4 in Section 2.3), we ensure comprehensive representation across all domains when dividing the training and test sets derived from these source tables. The distribution of source tables utilized for testing (to establish StatQA for the evaluation) and for training (to develop $\mathbb{D}_{\text{train}}$ for fine-tuning) is detailed in Table 5.
Table 5: Domains Distribution of Source Tables
Tables’ Domain StatQA (for test) $\mathbb{D}_{\text{train}}$ (for training)
Education 9 5
Medicine 4 7
Science 4 3
Engineering 3 4
Economy 10 13
Life 7 9
@@ C.3 Prompts
Appendix D More Experimental Results
To accommodate users only accessible to limited computational resources, or those with sufficient resources but are concerned about data privacy, a more extensive evaluation of open-source LLMs is essential. Accordingly, to provide references for these two groups of users, we conduct additional experiments on both smaller and larger representative open-source models, including Qwen2-0.5B-Instruct, Qwen2-1.5B-Instruct, Qwen2-72B-Instruct, Meta-Llama-3-8B, and Yi-34B-Chat. Table 10 illustrates experimental results for additional evaluation on open-source LLMs.
Qwen2-0.5B-Instruct is entirely inadequate for our task, presenting only 2.41% 1-shot overall accuracy and extremely low capabilities in all tasks. Qwen2-1.5B-Instruct demonstrates performance levels comparable to certain LLaMA models, specifically falling between LLaMA-2-13b and LLaMA-3-8b. Given that Qwen2-1.5B-Instruct is significantly smaller in parameters, although its performance is still not robust, it represents noteworthy and encouraging results for resource-constrained users.
In contrast, larger open-source LLMs exhibit significant enhancements, underscoring the importance of model scale. Remarkably, Qwen2-72B-Instruct demonstrates impressive performance with a 47.12% one-shot overall accuracy, surpassing all other open-source LLMs and being comparable to GPT models. Its strong capabilities are particularly promising for users with ample resources who prioritize data privacy and seek to minimize the reliance on proprietary solutions.
Table 10: More Open-source LLMs Experimental Results of $Acc(\mathbb{C},\mathbb{M})$ (%) on mini-StatQA. The 1st, 2nd, 3rd place results in all additional experiments here are highlighted in red, blue, and green respectively. The bold results are the best in each section.
Model Strategy Overall CA CTT DCT VT DS
Smaller Open-source LLMs
Qwen2 0.5B Instruct 0-shot 1.38 0.00 0.00 1.27 0.00 5.02
1-shot 2.41 0.00 0.00 3.81 1.23 6.18
Qwen2 1.5B Instruct 0-shot 15.99 0.00 1.56 1.27 18.03 52.12
1-shot 20.38 0.60 7.42 4.24 22.54 58.69
Larger Open-source LLMs
Qwen2 72B Instruct 0-shot 44.71 57.74 30.47 36.02 10.66 90.35
1-shot 47.12 66.07 30.47 28.81 21.72 91.89
LLaMA-3 70B 0-shot 27.09 35.12 2.34 11.02 19.67 67.95
1-shot 34.14 50.60 9.38 14.83 24.59 74.52
Yi 34B Chat 0-shot 28.03 41.07 13.67 14.41 0.00 72.59
1-shot 33.96 45.24 19.14 24.15 10.25 72.59
Appendix E More Error Analysis
E.1 Examples of Error Taxonomy